A wound coil that has an inductance of 180mH and a resistance of 35 is connected to a 100V 50Hz supply calculate the impedance of the coil the current the power factor and the apparent power consumed

1
Expert's answer
2022-03-15T10:46:28-0400

"R=35\Omega" ,"L=180mH,V=100V,f=50Hz"

a)Impedance (Z) of the coil

"R=35\Omega"

"X_L=2\pi fL=2\pi ×50×0.18=56.6\Omega"

"Z=\sqrt{R^2+X_L^2}=\sqrt{35^2+56.6^2}=66.5\Omega"


b)Current (I) consumed by the coil

"V=I\u00d7Z"

"\therefore I=\frac{V}{Z}=\frac{100}{66.5}=1.5A"


c)The power factor and phase angle,"\Phi"

"Cos\phi" "=\frac{R}{Z}or Sin\phi =\frac{X_L}{Z} or tan\phi =\frac{X_L}{R}"

"\therefore Cos\phi =\frac{R}{Z}=\frac{35}{66.5}=0.5263"

"Cos^{-1}(0.5263)=58.2\u00b0"


d)Apparent power (S) consumed by the coil

"P=V×I cos\phi =100×1.5cos(58.2°)=79W"

"Q=V×Isin\phi =100×1.5×sin(58.2°)=127.5VAr"

"S=V\u00d7I=100\u00d71.5=150VA"